OpenStudy (anonymous):

If sin(x) = 5/13, and x is in quadrant 1, then tan(x/2) equals what?

OpenStudy (anonymous):

sin^-1 (5/13)=22.61986495 tan(22.61986595/2)=0.2 or 1/5

OpenStudy (anonymous):

Use Pythagros therom 1.) sin x=5/13 sin is opp/hyp 2.) make triangle and solve for tangent tan=opp/adj 3.) so tanx= 5/12 Understand?

OpenStudy (anonymous):

Thanks guys!

OpenStudy (anonymous):

welcome, keep it simple, do it the way i did, you only use arc trig fucntions for angles